Which of the following is the slope between the two points (1, 5) and (6, 4)?

1/5
-1/5
5
-5

Answer by jim_thompson5910(35256)

m = (y2-y1)/(x2-x1)


m = ( 4- 5)/(6-1)


m = (-1)/5


m = -1/5


So the slope of the line that passes through the two points is -1/5


This means that the answer is choice B